OpenStudy (anonymous):

1/6 + ( ? ) = 11/18 Please help.

OpenStudy (anonymous):

11/18 - 1/6 = 11-3 / 18 = 8/18 = 4/9

OpenStudy (anonymous):

11/18-1/6=? 11/18-3/18=? 8/18=? ?=4/9